A company currently produces an average of 66 lamps a day. The material cost in the lamp production is $10 per lamp. The company’s utility cost is 16 dollars per day. It currently has 9 employees, who work 8 hours a day and get paid $15 per hour per employee. Please calculate the company's total productivity (unit/$).
Answer #1

Total Number of Lamps produced per day = 66

Material cost = $10 per lamp

= 10*66

= $660

Utility Cost = $16 per day

Total Number of employees = 9

Each employee work 8 hours per day

Total Work hours = 9*8

= 72 hours

Compensation per hour per employee = $15

Total compensation = 15*72

= $1080

Total Expense = (660+16+1080)

= $1756

Productivity = Total number of Lamps/Total Expense

= 66/1756

= 0.038 Lamps/$